### v1

# The scalar-torsion gravity corrections in the first-order inflationary models

Abstract
Abstract: In this work, we consider the corrections to the cosmological models based on the teleparallel equivalent of general relativity induced by the non-minimal coupling between scalar field and torsion. To determine these corrections, we consider a power-law parametrization of the deviations between the teleparallel equivalent of general relativity and the scalar-torsion gravity. The estimates of the possible influence of the non-minimal coupling of the scalar field and torsion on the background cosmological parameters and the parameters of cosmological perturbations for the inflationary models implying linear relation between tensor-to-scalar ratio and spectral index of the scalar perturbations are obtained. A procedure for verifying these inflationary models based on observational constraints on the values of cosmological perturbation parameters is also considered.
I Introduction
In modern cosmological models, an important role is played by the description of two stages of the accelerated expansion of the universe, namely the primary inflation of the early universe [1, 2] and the second inflation in the modern era [3] of the evolution of the universe. To describe the cosmological dynamics of the universe at the different stages of its evolution and observed large-scale structure of the universe, general relativity [1, 2] and various modified theories of gravity are considered [4, 5, 6, 7]. One such approach is the use of the teleparallel equivalent of general relativity (TEGR) to describe gravity, where torsion scalar $T$ is considered instead of the scalar curvature $R$ [8, 9]. Also, the different modifications of TEGR are considered to construct and analyze the cosmological models  [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24].
Equations of cosmological dynamics for the case of $f(T)$-gravity is the second order as in the case of general relativity, however, instead of using the curvature defined through the Levi-Civita connection, in the case of $f(T)$-gravity the Weitzenböck connection is used, which implies zero curvature and non-zero torsion [25, 26, 27]. The fact that the dynamic equations for $f(T)$-gravity are second order [25] makes these theories simpler than those modified by curvature invariants (see, for example, in [4, 5, 6, 7]), and a study of this type of model is of considerable interest for theoretical cosmology. It should also be noted that the speed of gravitational waves for the case of $f(T)$-gravity is equal to the speed of light in vacuum, which is similar to both the case of general relativity and scalar-tensor theories of gravity and satisfies the restrictions on the speed of gravitational waves from astrophysical objects according to modern observations [28].
Thus, the teleparallel equivalent of general relativity (TEGR) and its modifications are the basis for the construction of cosmological models which can be verifiable by observational data, that corresponds to the growing interest in this type of gravity models and their physical applications in modern cosmology  [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. When constructing and analyzing cosmological models based on TEGR modifications, the following questions arises: how do these modifications affect cosmological parameters and what possible physical effects do they correspond to? To answer these questions, the approach associated with a model-independent assessment of the influence of modified theories of gravity on cosmological parameters seems promising. In  [29, 30, 31, 32, 33, 34, 35] methods were proposed for analyzing cosmological models with modified gravity theories (scalar-tensor gravity and Einstein-Gauss-Bonnet gravity) based on functional and parametric connections to the Einstein gravity in the Friedman-Robertson-Walker space-time. This approach made it possible to directly assess the influence of these modifications of the Einstein gravity on the background dynamics and parameters of cosmological perturbations. In  [36] this approach was used for the reconstruction of the scalar-torsion gravity theories from the physical potentials of a scalar field.
In this work, we present the model-independent approach to estimate the influence of non-minimal coupling between a scalar field and torsion on the cosmological parameters. In the context of the proposed approach to this interpretation of the description of the inflationary scenario, we consider the influence of the non-minimal coupling of the scalar field and torsion on the cosmological dynamics, the scalar field potential and the parameters of cosmological perturbations. The proposed power-law relationship between the coupling function and the Hubble parameter allows us to reduce the assessment of such influence to the estimation of the possible values of one constant parameter. It was also shown, that the corrections inspired by non-minimal coupling of the scalar field and torsion lead to the verification of the inflationary models implying a linear relation between tensor-to-scalar ratio and spectral index of the scalar perturbations by the observational constraints in contrast to the case of TEGR.
The paper is organized as follows. In Section II we consider dynamical equations for cosmological models based on the scalar-torsion gravity. Section III proposes a power-law relation between the non-minimal coupling function and the Hubble parameter $F=(H/\lambda)^{n}$. The implications of this relation are also discussed. In Section IV, we obtain the parameter $\lambda$ in explicit form based on the conditions for the correct completion of the inflationary stage. Section V provides restrictions on the values of parameter $n$ and on the possible types of inflationary models. In Section VI it was shown that the considered inflationary models with non-minimal coupling between scalar field and torsion can be verified from observational data, unlike models with minimal coupling. Estimates of the propagation velocity of scalar perturbations for the verified inflationary models were also given. A summary of our investigations is presented in Sec. VII.
II Dynamical equations for cosmological models based on the scalar-torsion gravity
The action of generalized scalar-torsion theory with self-interaction of Galilean-like field in the system of units $M^{2}_{p}=(8\pi G)^{-1}=c=1$ can be defined as follows [18]
$S=\int d^{4}xe\left[\frac{1}{2}F(\phi)T+P(\phi,X)-\square\phi G(\phi,X)\right],$ (1)
where $F(\phi)$ is the function of the scalar field $\phi$, $P$, $G$ are functions of the scalar field and its kinetic energy $X=-\frac{1}{2}\phi_{,\mu}\phi^{,\mu}$ and $\phi_{,\mu}=\partial_{\mu}\phi=\frac{\partial\phi}{\partial x^{\mu}}$, $T$ is the torsion scalar, $e^{A}_{\mu}$ are the tetrad components in a coordinate basis, $e=det(e^{A}_{\mu})=\sqrt{-g}$ and $G(\phi,X)$ is the Galileon-type field self-interaction. Also, the function $P$ can be considered as the function of the potential of a scalar field and its kinetic energy $P=P\left(V(\phi),X\right)$.
In this paper, we consider a special case of negligible self-interaction of a scalar field
$\frac{1}{2}F(\phi)T\gg\square\phi G(\phi,X).$ (2)
In this case, the action Eq. (1) is reduced to the following form
$S\approx\int d^{4}xe\left[\frac{1}{2}F(\phi)T+P(\phi,X)\right].$ (3)
Thus, we will consider the nonlinear self-interaction of the scalar field as a higher-order effect compared to the non-minimal coupling of the scalar field and torsion.
The tetrad field
$e^{A}_{\mu}=diag\left(1,a,a,a\right),$ (4)
corresponds to the homogeneous and isotropic Friedman-Robertson-Walker (FRW) universe with metric
$ds^{2}=-dt^{2}+a^{2}\delta_{ij}dx^{i}dx^{j},$ (5)
where $a=a(t)$ is the scale factor which dependent on cosmic time $t$ for expanding universe.
Thus, the background equations for the action Eq. (3) and tetrad field (4) are [18, 19, 20]
$P-2XP_{,X}-3H^{2}F=0,$ (6)
$P-3H^{2}F-2\dot{H}F-2\dot{F}H=0,$ (7)
$P_{,\phi}-2XP_{,\phi X}-3H^{2}_{,\phi}-3HP_{,X}\dot{\phi}$
$-\left[P_{,X}+2P_{,XX}\right]\ddot{\phi}=0,$ (8)
where overdot means the derivative with respect to cosmic time.
Let us consider a particular case where the function $P=P(\phi,X)$ is defined as follows
$P=-\omega X+V(\phi)=-\frac{\omega}{2}\dot{\phi}^{2}+V(\phi),$ (9)
where $\omega$ is a some constant.
Thus, from Eqs. (7)–(9) we obtain
$V(\phi)=3H^{2}F+\dot{H}F+\dot{F}H,$ (10)
$\omega\dot{\phi}^{2}=-2H\dot{F}-2F\dot{H}.$ (11)
For the case of minimal coupling $F=1$, these models are reduced to the case of the inflationary models based on the teleparallel equivalent of general relativity.
III Power-law parametrization of coupling function
In order to estimate the influence of the non-minimal coupling on the cosmological parameters, we consider some connection between coupling function $F=F(\phi)$ and the Hubble parameter $H=H(t)$ to parameterize the influence of the non-minimal coupling between the scalar field and torsion on the parameters of cosmological models under consideration.
We define this connection as the power-law one, namely
$F(\phi(t))=\left(\frac{H(t)}{\lambda}\right)^{n},$ (12)
where $\lambda$ is a positive constant and $n$ is the parameter that defines the influence of the non-minimal coupling between the scalar field and torsion. For $n=0$ one has the case $F=1$ corresponding to minimal coupling or TEGR.
Also, we note, that relation (12) can be defined in parametric form for any solutions of dynamic equations (10)–(11), and for the special case $n=0$, we obtain the minimal coupling for any Hubble parameter.
Dynamic equations (10)–(11), in this case, have the following form
$V(\phi)=3\lambda^{-n}H^{2+n}+\lambda^{-n}\dot{H}H^{n}(1+n),$ (13)
$\omega\dot{\phi}^{2}=-2\lambda^{-n}\dot{H}H^{n}(n+1).$ (14)
In this work, unlike paper [36], where kinetic function was considered as $\omega=\omega(\phi)$, we will consider $\omega=const$ corresponding to the canonical (for $\omega>0$) and phantom (for $\omega<0$) scalar fields.
Thus, equations (13)–(14) can be rewritten as
$V(\phi)=\frac{H^{n}}{\lambda^{n}}\left(3H^{2}+\dot{H}(1+n)\right)$
$=F(\phi)\left(3H^{2}+\dot{H}(1+n)\right),$ (15)
$\omega\dot{\phi}^{2}=-2\left(\frac{H}{\lambda}\right)^{n}\dot{H}(n+1)=-\frac{2}{\lambda^{n}}\frac{d}{dt}\left(H^{n+1}\right),$ (16)
where we will consider
$\omega=\lambda^{-n}=const,$ (17)
as a convenient normalization of the constant $\omega$ implying canonical scalar field $\omega=1$ for the minimal coupling $n=0$.
Thus, for the case of the minimal coupling $n=0$, dynamic equations (III)–(16) are reduced to the well-known expressions [2]
$\tilde{V}(\phi)=3h^{2}+\dot{h}=3h^{2}-\frac{1}{2}\dot{\phi}^{2},$ (18)
$\dot{\phi}^{2}=-2\dot{h},$ (19)
where $h=h(t)>0$ is the Hubble parameter corresponding to the minimal coupling.
The quasi de-Sitter expansion of the early universe $h\approx const$ implies the following slow-roll condition $X=\frac{1}{2}\dot{\phi}^{2}\ll V$ at the inflationary stage.
Let us introduce the connection between $H$ and $h$ like follow
$H=h^{\frac{1}{1+n}},$ (20)
$\dot{H}=\left(\frac{1}{1+n}\right)\dot{h}h^{-\frac{n}{1+n}},$ (21)
we can rewrite equations (III)–(19) as follows
$V(\phi)=\lambda^{-n}\left[3h^{\frac{2+n}{1+n}}+\dot{h}\right]=\lambda^{-n}\left[3h^{\frac{2+n}{1+n}}+\frac{1}{2}\dot{\phi}^{2}\right],$ (22)
$\dot{\phi}^{2}=-2\dot{h}.$ (23)
Therefore, the kinetic energy of the scalar field for non-minimal and minimal coupling is the same $X=\tilde{X}=\frac{1}{2}\dot{\phi}^{2}$ for relation (20). This relation allows us to consider the fulfillment of the inflationary slow-roll conditions $\tilde{X}\ll\tilde{V}$ and $X\ll V$ for these two cases at once. Also, relation (20) implies the same evolution of the scalar field for both these cases $\phi(t)=\tilde{\phi}(t)$ up to the choice of a constant.
Also, using Ivanov-Salopek-Bond approach [2] based on the relation $\dot{h}=\dot{\phi}h_{,\phi}=\dot{\phi}\frac{dh(\phi)}{d\phi}$, we can rewrite (22)–(23) as
$V(\phi)=\lambda^{-n}\left\{3\left[h(\phi)\right]^{\frac{2+n}{1+n}}-2(h_{,\phi})^{2}\right\},$ (24)
$\dot{\phi}=-2h_{,\phi}.$ (25)
Also, we note, that under condition $01+n>$ 0 for $\dot{h}<0$ one has $\dot{H}<0$ as well. Thus, under condition $01+n>$ 0 the inflationary stage is being implemented for the case of non-minimal coupling if inflation occurs for the case of minimal coupling.
The condition of inflationary stage $01+n>$ 0 restricts the possible values of the characterization constant as follows
$-1<n.$ (26)
The coupling function (12) in terms of the Hubble parameter for minimal coupling is
$F(\phi)=\lambda^{-n}\left[h(\phi)\right]^{\frac{n}{1+n}}.$ (27)
Also, under the slow-roll conditions $\tilde{X}\ll\tilde{V}$ and $X\ll V$, from expressions (III), (18) and (24) we get
$\tilde{V}(\phi)\simeq 3h^{2},$ (28)
$V(\phi)\simeq 3\lambda^{-n}H^{2+n}=3\lambda^{-n}h^{\frac{2+n}{1+n}}.$ (29)
Thus, from (28)–(29) we obtain the following relation between potentials for non-minimal and minimal coupling
$V(\phi)\simeq 3\lambda^{-n}\left(\frac{\tilde{V}(\phi)}{3}\right)^{\frac{2+n}{2(1+n)}}.$ (30)
Also, from expressions (27) and (28) we obtain the relation between the coupling function and the potential for minimal coupling
$F(\phi)\simeq\lambda^{-n}\left(\frac{\tilde{V}(\phi)}{3}\right)^{\frac{n}{2(1+n)}}.$ (31)
Further, based on relations (20)–(21), we will express the slow-roll parameters for the case of non-minimal coupling in terms of the Hubble parameter $h=h(t)$ as follows
$\epsilon=-\frac{\dot{H}}{H^{2}}=-\frac{\dot{h}}{(1+n)}h^{-\frac{2+n}{1+n}},$ (32)
$\delta=-\frac{\ddot{H}}{2H\dot{H}}=\frac{n\dot{h}}{2(1+n)}h^{-\frac{2+n}{1+n}}-\frac{\ddot{h}}{2\dot{h}}h^{-\frac{1}{1+n}}.$ (33)
For the case $n=0$, from (32)–(33) one has the slow-roll parameters for inflation with minimal coupling [1, 2]
$\tilde{\epsilon}=-\frac{\dot{h}}{h^{2}}=2\left(\frac{h_{,\phi}}{h}\right)^{2},$ (34)
$\tilde{\delta}=-\frac{\ddot{h}}{2\dot{h}h}=2\frac{h_{,\phi\phi}}{h}.$ (35)
Thus, to estimate the corrections induced by the non-minimal coupling between scalar field and torsion, it is necessary to determine two constant parameters: $\lambda$ and $n$. We will determine the constant $\lambda$ based on the conditions for the completion of the inflationary stage. Also, we will determine the lower limit on the parametrization constant $n$ based on the analysis of the cosmological perturbations parameters.
IV State parameter at the end of the inflationary stage
Now, we determine the constant parameter $\lambda$ from the conditions on the state parameter at the end of the stage of cosmological inflation. In cosmological models with minimal coupling, the state parameter of a scalar field is determinate as follows [1, 2]
$\tilde{w}=\frac{\tilde{p}}{\tilde{\rho}}=\frac{\tilde{X}-\tilde{V}}{\tilde{X}+\tilde{V}}=-1-\frac{2}{3}\frac{\dot{h}}{h^{2}}=-1+\frac{2}{3}\tilde{\epsilon}.$ (36)
where $\tilde{p}$ and $\tilde{\rho}$ are the pressure and energy density of a scalar field correspondingly.
At the inflationary stage, due to condition $\tilde{\epsilon}\ll 1$ from expressions (36) one has $\tilde{w}\simeq-1$. The end of the accelerated expansion stage of the early universe is defined by condition $\tilde{\epsilon}=1$ with corresponding state parameter $\tilde{w}_{E}=-1/3$ which follows from (36) for the case of the minimal coupling [1, 2].
For the case of non-minimal coupling $n\neq 0$, the state parameter is
$w=\frac{p}{\rho}=\frac{X-V}{X+V}=\frac{-3\lambda^{-n}h^{\frac{2+n}{1+n}}-\dot{h}(1+\lambda^{-n})}{3\lambda^{-n}h^{\frac{2+n}{1+n}}-\dot{h}(1-\lambda^{-n})}.$ (37)
Taking to account expression (32), from expression (37) we obtain
$w=\frac{-3+(\lambda^{n}+1)(1+n)\epsilon}{3+(\lambda^{n}-1)(1+n)\epsilon}.$ (38)
At the inflationary stage, due to conditions $\tilde{\epsilon}\ll 1$ and $\epsilon\ll 1$ from expressions (36) and (38) one has $\tilde{w}\simeq-1$ and $w\simeq-1$.
For the special case $n=2$, due to expression (38), at the end of inflation $\epsilon=1$ the state parameter is $w_{E}=1$ for any value of the constant parameter $\lambda$. This case corresponds to the post-inflationary stage of kinetic energy domination (kination), which precedes the further stage of radiation domination [37, 38, 39].
For the case $n\neq 2$, at the end of inflation $\epsilon=1$ from condition $w_{E}=-1/3$ and expression (38), we obtain following relation
$\lambda^{n}=\frac{2-n}{2(1+n)},$ (39)
corresponding to the standard completion of the inflationary stage.
Also, from condition $\lambda^{n}>0$, we obtain following restrictions on the parametrization constant $-1<n<2$ for this case.
Thus, from equations (30)–(31) and (39), we obtain following relations
$V(\phi)\simeq\frac{6(1+n)}{(2-n)}\left(\frac{\tilde{V}(\phi)}{3}\right)^{\frac{2+n}{2(1+n)}},$ (40)
$F(\phi)\simeq\frac{2(1+n)}{(2-n)}\left(\frac{\tilde{V}(\phi)}{3}\right)^{\frac{n}{2(1+n)}},$ (41)
which characterize the influence of the non-minimal coupling on the potential and define the coupling function.
To determine what type of completion of the inflationary stage is implemented for the models under consideration, it is necessary to determine the upper limit on the parametrization constant $n$ in addition to the lower limit (26). This estimate can be obtained on the base of the analysis of the influence of non-minimal coupling between a scalar field and torsion on the parameters of the cosmological perturbations in the inflationary models under consideration.
V Adjusted parameters of cosmological perturbations
From the standpoint of the inflationary paradigm, the origin of the large-scale structure of the universe is due to quantum fluctuations of the scalar field and corresponding scalar perturbations of the space-time metric at the inflationary stage of the evolution of the universe. Also, the theory of cosmological perturbations predicts the presence of relict gravitational waves or tensor perturbations in the present era which were also induced at the inflationary stage [1, 2].
The influence of cosmological perturbations on the anisotropy and polarization of CMB makes it possible to determine the correctness of inflationary models based on observational constraints on the values of the parameters of cosmological perturbations [1, 2]. According to modern observational data [40, 41], constraints on the parameters of cosmological perturbations are
$P_{S}=2.1\times 10^{-9},$ (42)
$n_{S}=0.9663\pm 0.0041,$ (43)
$r<0.032,$ (44)
where $P_{S}$ is the power spectrum of scalar perturbations, $n_{S}$ is the spectral index of scalar perturbations and $r$ is the tensor-to-scalar ratio.
It is obvious that the non-minimal coupling of the scalar field and torsion affects the evolution of cosmological perturbations and induces some corrections to the values of the parameters of cosmological perturbations relative to the case of minimal coupling. The analysis of the evolution of cosmological disturbances in inflationary models based on the TEGR and its modifications was considered earlier in Refs. [15, 16, 17, 18, 19].
In Ref. [18] cosmological perturbations for the inflationary models based on the generalized scalar-torsion gravity were considered based on the Arnowitt-Deser-Misner (ADM) decomposition of the tetrad field to define the perturbed metrics. By this approach, the parameters of cosmological perturbations at the crossing of the Hubble radius $k=aH$ can be written as follows [18]
${\mathcal{P}}_{S}=\frac{H^{2}}{8\pi^{2}Q_{S}c_{S}^{3}}=\frac{H^{2}}{8\pi^{2}Fc_{S}\epsilon_{s}},$ (46)
${\mathcal{P}}_{T}=\frac{H^{2}}{2\pi^{2}Q_{T}}=\frac{2H^{2}}{\pi^{2}F},$
$n_{S}-1=-2\epsilon-\delta_{F}-\eta_{s}-s,$ (49)
$n_{T}=-2\epsilon-\delta_{F},$
$r=\frac{{\mathcal{P}}_{T}}{{\mathcal{P}}_{S}}=16c_{s}\epsilon_{s},$
where the stability parameters are
$Q_{S}=\frac{w_{1}}{3w_{2}^{2}}\left(9w_{2}^{2}+4w_{1}w_{3}\right),~{}~{}~{}~{}Q_{T}=\frac{1}{4}F,$ (50)
the slow-roll parameters are
$\epsilon=-\frac{\dot{H}}{H^{2}},~{}~{}~{}~{}\delta=-\frac{\ddot{H}}{2H\dot{H}},~{}~{}~{}~{}\delta_{F}=\frac{\dot{F}}{HF},$ (51)
$\epsilon_{s}=\frac{Q_{S}c_{S}^{2}}{F},~{}~{}~{}~{}\eta_{s}=\frac{\dot{\epsilon}_{s}}{H\epsilon_{s}},~{}~{}~{}~{}s=\frac{\dot{c}_{S}}{Hc_{S}},$ (52)
and the velocity of the scalar perturbations is
$c_{S}^{2}=\frac{3\left(2Hw_{1}w_{2}+4\dot{w_{1}}w_{2}-2w_{1}\dot{w_{2}}-w_{2}^{2}\right)}{9w_{2}^{2}+4w_{1}w_{3}}.$ (53)
Also, we note, that all slow-roll parameters satisfy the condition $\epsilon,\delta,\delta_{F},\epsilon_{s},\eta_{s},s\ll 1$ during inflation, and condition $\dot{H}<0$ implies $\epsilon>0$ as well.
For the inflationary models based on the action (1) functions $w_{1}$, $w_{2}$, $w_{3}$ can be defined as [18]
$w_{1}=F,~{}~{}~{}~{}w_{2}=2HF,$ (54)
$w_{3}=-3XP_{,X}-9H^{2}F=\frac{3}{2}\omega\dot{\phi}^{2}-9H^{2}F$
$=-3F\dot{H}-3H\dot{F}-9H^{2}F,$ (55)
and the velocity of the tensor perturbations is equal to the speed of light in vacuum $c_{T}^{2}=1$.
As one can see, in the general case, the velocity of propagation of scalar perturbations is a function of cosmic time $c_{S}=c_{S}(t)$. Thus, in the general case, corrections induced by the non-minimal coupling of the scalar field and torsion must be considered for each specific model separately. Nevertheless, for the power-law connection between the non-minimal coupling function and the Hubble parameter $F=(H/\lambda)^{n}$ from (53)–(V), we obtain the following expression for the velocity of propagation of scalar perturbations
$c_{S}^{2}=\frac{1-n}{1+n}=const,$ (56)
for an arbitrary model’s parameters.
The condition $c_{S}^{2}>0$ for expression (56) leads to the following conditions on the parametrization constant
$-1<n<1,$ (57)
which correspond to lower limit (26).
Thus, we have three possible regimes of scalar perturbations propagation, namely
  1. 1.
    sub-luminal regime with $10<c_{S}<$ 1 for $10<n<$ 1,
  2. 2.
    luminal regime with $c_{S}=1$ for $n=0$,
  3. 3.
    super-luminal regime with $c_{S}>1$ for $-1<n<0$.
A super-luminal propagation cosmological perturbations is possible in the presence of the time-dependent homogeneous scalar field [42, 43, 44] and for different modified gravity theories [45, 46, 47] as well. Also, we note, that due to constraint (57), the possible post-inflationary kination stage (for $n=2$) can not be realized in these cosmological models.
Figure 1: The dependence of the velocity of scalar perturbations from the value of the parametrization constant $n$.
The difference in the propagation velocities of scalar and tensor perturbations corresponds to the fact that they cross the Hubble radius with different wavelengths, namely
$c_{S}k_{S}=aH,~{}~{}~{}~{}k_{S}=\frac{2\pi a}{\lambda_{S}},$ (58)
$c_{T}k_{T}=aH,~{}~{}~{}k_{T}=\frac{2\pi a}{\lambda_{T}},$ (59)
$\frac{c_{S}}{c_{T}}=\frac{k_{T}}{k_{S}}=\frac{\lambda_{S}}{\lambda_{T}}=\sqrt{\frac{1-n}{1+n}},$ (60)
unlike the case of the minimal coupling $n=0$ for which $c_{S}=c_{T}=1$ and the wavelengths of scalar and tensor perturbations are equal $\lambda_{S}=\lambda_{T}$.
Also, for the case $F=(H/\lambda)^{n}$, from (50) we obtain
$Q_{S}=-(1+n)\frac{\dot{H}}{H^{2}}\left(\frac{H}{\lambda}\right)^{n}$
$=(1+n)\left(\frac{H}{\lambda}\right)^{n}\epsilon>0,$ (61)
$Q_{T}=\frac{1}{4}F=\frac{1}{4}\left(\frac{H}{\lambda}\right)^{n}>0.$ (62)
The conditions $Q_{S}>0$ and $Q_{T}>0$ mean that there are no Laplacian instabilities and ghosts in these models [18].
The slow-roll parameters (51)–(52) for the case $F=(H/\lambda)^{n}$ are related as follows
$\epsilon_{s}=(1-n)\epsilon,~{}~{}~{}~{}~{}~{}\delta_{F}=-n\epsilon,$
$\eta_{s}=-2\delta+2\epsilon,~{}~{}~{}~{}~{}~{}s=0.$ (63)
Thus, parameters of cosmological perturbations (46)–(49) can be written as
${\mathcal{P}}_{S}=\frac{(2-n)H^{2-n}_{\ast}}{8\pi^{2}(1-n^{2})\sqrt{\frac{1-n}{1+n}}\,\epsilon_{\ast}},$ (64)
${\mathcal{P}}_{T}=\frac{2\lambda^{n}}{\pi^{2}}H^{2-n}_{\ast},$ (65)
$n_{S}-1=(n-4)\epsilon_{\ast}+2\delta_{\ast},$ (66)
$r=16(1-n)\sqrt{\frac{1-n}{1+n}}\,\epsilon_{\ast},$ (67)
$n_{T}=-(2-n)\epsilon_{\ast}.$ (68)
For the case of the minimal coupling $n=0$ we obtain
$\tilde{{\mathcal{P}}}_{S}=\frac{h^{2}_{\ast}}{8\pi^{2}\,\tilde{\epsilon}_{\ast}},$ (69)
$\tilde{{\mathcal{P}}}_{T}=\frac{2h^{2}_{\ast}}{\pi^{2}},$ (70)
$\tilde{n}_{S}-1=-4\tilde{\epsilon}_{\ast}+2\tilde{\delta}_{\ast},$ (71)
$\tilde{r}=16\tilde{\epsilon}_{\ast},$ (72)
$\tilde{n}_{T}=-2\tilde{\epsilon}_{\ast}.$ (73)
We note, that all parameters of cosmological perturbations are considered at the crossing of the Hubble radius ($k=aH$ and $\tilde{k}=\tilde{a}h$ for some cosmic times $\tilde{t}=\tilde{t}_{\ast}$ and $t=t_{\ast}$, where $t_{\ast}\neq\tilde{t}_{\ast}$ and, correspondingly, $H^{n+1}_{\ast}\neq h_{\ast}$).
Since in both cases (non-minimal and minimal coupling) cosmological models can satisfy condition (42), the relation between the cosmic times of the crossing of the Hubble radius $\xi=\left(t_{\ast}/\tilde{t}_{\ast}\right)$ can be obtained from the equation for certain inflationary models
${\mathcal{P}}_{S}=\tilde{{\mathcal{P}}}_{S}=2.1\times 10^{-9},$ (74)
where $\xi=1$ for the case $n=0$.
Since condition (42) can always be satisfied by choosing the time of crossing the Hubble radius or model’s parameters, we consider the spectral index of scalar perturbations and the tensor-scalar ratio as the main parameters necessary for verifying inflationary models by observational constraints (43)–(44). Also, in accordance with restrictions (57), condition $\epsilon_{\ast}>0$ and expression (68), on has red tilt tensor spectrum $n_{T}<0$ for arbitrary cosmological models under consideration.
Also, we note, that expressions for the slow-roll parameters (32)–(35) leads to the following relation
$\frac{\delta}{\epsilon}=(1+n)\frac{\tilde{\delta}}{\tilde{\epsilon}}-\frac{n}{2}.$ (75)
Since relation (75) must be satisfied for any cosmic times, we get following conditions
$\tilde{\delta}/\tilde{\epsilon}=const,$ (76)
$\delta/\epsilon=const,$ (77)
which restrict the possible types of inflationary models.
VI Inflationary models with linear dependence $r\sim(1-n_{S})$
Now we will consider the dependence of the tensor-scalar ratio on the spectral index of scalar perturbations $r=r(1-n_{S})$. Since the value of the spectral index of scalar perturbations is $n_{S}\simeq 0.97$ and $1-n_{S}\simeq 0.03\ll 1$, we can write the dependence $r=r(1-n_{S})$ as follows
$r=\sum^{\infty}_{k=0}\beta_{k}(1-n_{S})^{k}=\beta_{0}+\beta_{1}(1-n_{S})+$
$\beta_{2}(1-n_{S})^{2}+...,$ (78)
where $(1-n_{S})\ll 1$ is the small parameter of expansion and $\beta_{k}$ are the constant coefficients.
Since, the zeroth order term in this expansion $r(0)=\beta_{0}=0$ from condition $r(n_{S}=1)=0$ corresponding to the flat Harrison-Zel’dovich spectrum [48, 49], we can rewrite expression (VI) in the following form
$r=\sum^{\infty}_{k=1}\beta_{k}(1-n_{S})^{k}=\beta_{1}(1-n_{S})$
$+\beta_{2}(1-n_{S})^{2}+...$ (79)
The first term in expression (VI) makes the main contribution to the value of the tensor-scalar ratio. Thus, we can consider the dependence $r=r(1-n_{S})$ with a reasonable degree of accuracy at the first order
$r\simeq\beta_{1}(1-n_{S}),$ (80)
to analyze the influence of non-minimal coupling on the parameters of inflationary models.
To obtain this dependence, we consider the inflationary models with minimal coupling and constant relation between slow-roll parameters during inflation
$\frac{\tilde{\delta}}{\tilde{\epsilon}}=\frac{\tilde{\delta}_{\ast}}{\tilde{\epsilon}_{\ast}}=const.$ (81)
After substituting relation (81) into (71)–(72), we obtain a linear relationship between the tensor-scalar ratio and the spectral index of scalar perturbations
$\tilde{r}=\tilde{\beta}_{1}(1-\tilde{n}_{S}).$ (82)
Also, based on expressions (19), (28) and (34)–(35), we can express the slow-roll parameters in terms of potential in the slow-roll approximation [2]
$\tilde{\epsilon}=-\frac{\dot{h}}{h^{2}}=2\left(\frac{h_{,\phi}}{h}\right)^{2}\simeq\frac{1}{2}\left(\frac{\tilde{V}_{,\phi}}{\tilde{V}}\right)^{2},$ (83)
$\tilde{\delta}=-\frac{\ddot{h}}{2\dot{h}h}=2\frac{h_{,\phi\phi}}{h}\simeq\frac{\tilde{V}_{,\phi\phi}}{\tilde{V}}-\frac{1}{2}\left(\frac{\tilde{V}_{,\phi}}{\tilde{V}}\right)^{2},$ (84)
where we use the following relations
$\dot{h}=-2h^{2}_{,\phi},~{}~{}~{}~{}~{}\ddot{h}=8h^{2}_{,\phi}h_{,\phi\phi},~{}~{}~{}~{}~{}h\simeq\left(\frac{V}{3}\right)^{1/2}.$ (85)
Thus, from condition (81) and equations (83)–(84), we obtain two types of inflationary potential:
$\tilde{V}(\phi)\sim\phi^{p},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}{\text{f}or}~{}~{}~{}\tilde{\delta}/\tilde{\epsilon}\neq 1,$ (86)
$\tilde{V}(\phi)\sim\exp(-\alpha\phi),~{}~{}~{}~{}{\text{f}or}~{}~{}~{}\tilde{\delta}/\tilde{\epsilon}=1.$ (87)
where $p$ and $\alpha$ are a some constants.
Models of cosmological inflation based on these potentials have been considered previously in a large number of papers (see, for example, in [50, 51, 52, 53, 54, 55, 56]).
Also, from (29), (32)–(33) and (85), we obtain the following expressions for the slow-roll parameters for the case of non-minimal coupling
$\epsilon=\frac{2\left(\frac{d}{d\phi}h\left(\phi\right)\right)^{2}h\left(\phi\right)^{-\frac{2+n}{1+n}}}{1+n},$ (88)
$\delta=-\frac{2n\left(\frac{d}{d\phi}h\left(\phi\right)\right)^{2}h\left(\phi\right)^{-\frac{2+n}{1+n}}}{2+2n}$
$+2\left(\frac{d^{2}}{d\phi^{2}}h\!\left(\phi\right)\right)h\left(\phi\right)^{-\frac{1}{1+n}},$ (89)
$h\left(\phi\right)\simeq\left(\frac{\lambda^{n}}{3}V\left(\phi\right)\right)^{\frac{1+n}{2+n}}.$ (90)
Using these expressions, we will check the compliance with condition $\delta/\epsilon=const$ in the slow-roll approximation, since relation (75) was obtained for exact expressions of the slow-roll parameters.
VI.1 Inflationary models with power-law potential and minimal coupling
Now, we consider the inflationary model with power-law potential [50, 51, 52, 53, 54, 55, 56]
$\tilde{V}(\phi)=\tilde{V}_{0}\phi^{p},$ (91)
where $\tilde{V}_{0}$ and $p$ are a some constants.
From (83)–(84), we obtain
$\tilde{\epsilon}=\frac{p^{2}}{2\phi^{2}}>0,$ (92)
$\frac{\tilde{\delta}}{\tilde{\epsilon}}=\frac{\tilde{\delta}_{\ast}}{\tilde{\epsilon}_{\ast}}=\left(\frac{p-2}{p}\right)=const.$ (93)
Since the slow-roll parameters (92)–(93) contain an unknown arbitrary constant $p$, in this case, it is necessary to obtain an additional dependence of the slow-roll parameters on the e-folds number $\tilde{N}$.
Firstly, we define the value of the scalar field at the end of inflation $\phi=\phi_{e}$. From condition $\epsilon=1$ and expression (92) we obtain
$\phi^{2}_{e}=\frac{p^{2}}{2}.$ (94)
Secondly, we define the e-folds number between the crossing of the Hubble radius and the end of inflation [1, 2]
[…]
Further after substituting (91) into (VI.1), we obtain
$\tilde{N}\simeq\frac{1}{2p}\left(\phi^{2}_{\ast}-\phi^{2}_{e}\right)=\frac{1}{2p}\left(\phi^{2}_{\ast}-\frac{p^{2}}{2}\right).$ (96)
Thus from (96), we obtain the value of the scalar field at the crossing of the Hubble radius
$\phi^{2}_{\ast}=\frac{p}{2}\left(p+4\tilde{N}\right).$ (97)
Also, from (92)–(93) and (97) we can write the slow-roll parameters at the crossing of the Hubble radius as
$\tilde{\epsilon}_{\ast}=\frac{p}{p+4\tilde{N}},~{}~{}~{}~{}~{}\tilde{\delta}_{\ast}=\frac{p-2}{p+4\tilde{N}}.$ (98)
From (71)–(72) and (98) we obtain the following expressions for the spectral tilt of the scalar perturbations and tensor-to-scalar ratio
$1-\tilde{n}_{S}=\frac{2(p+2)}{p+4\tilde{N}},$ (99)
$\tilde{r}=\frac{16p}{p+4\tilde{N}}.$ (100)
Thus, from equations (99)–(100) we obtain following dependencies of the tensor-to-scalar ratio from the spectral index of the scalar perturbations
$\tilde{r}=\frac{8p}{p+2}(1-\tilde{n}_{S}),$ (101)
$\tilde{r}=\frac{16\left[(1-\tilde{n}_{S})\tilde{N}-1\right]}{2\tilde{N}-1}.$ (102)
Due to the observational value of the spectral tilt of the scalar perturbations $\tilde{n}_{S}=0.9663\pm 0.0041$ for a fairly wide range of the e-folds number $7050<\tilde{N}<$ 70, from (102) we obtain $0.200.08<\tilde{r}<$ 0.20. Thus, for the minimal coupling, these models don’t correspond to the observational constraint (44).
Also, from expression (99), we obtain:
$p_{min}(\tilde{n}_{S}=0.9704,\tilde{N}=50)\simeq 1,$ (103)
$p_{max}(\tilde{n}_{S}=0.9622,\tilde{N}=70)\simeq 3.34.$ (104)
Thus, in the slow-roll approximation, we can estimate the values of the parameter $p$ as $3.41<p<$ 3.4.
VI.2 Inflationary models with power-law potential and non-minimal coupling
Now, we estimate the influence of non-minimal coupling between the scalar field and torsion on the cosmological parameters. From equations (40)–(41) and (91) we obtain the following expressions for the potential and non-minimal coupling function
$V(\phi)\simeq V_{0}\phi^{k},$ (105)
$F(\phi)\simeq F_{0}\phi^{m},$ (106)
where
$V_{0}=\frac{6(1+n)}{(2-n)}\left(\frac{\tilde{V}_{0}}{3}\right)^{\frac{2+n}{2(1+n)}},~{}~{}~{}k=\frac{p(2+n)}{2(1+n)},$ (107)
$F_{0}=\frac{2(1+n)}{(2-n)}\left(\frac{\tilde{V}_{0}}{3}\right)^{\frac{n}{2(1+n)}},~{}~{}~{}m=\frac{pn}{2(1+n)}.$ (108)
After substituting (105) and (107) into (88)–(90) we obtain
$\frac{\delta}{\epsilon}=\frac{\delta_{\ast}}{\epsilon_{\ast}}=(1+n)\left(\frac{p-2}{p}\right)-\frac{n}{2}=const,$ (109)
that coincides with relation (75).
Also, from (66)–(67) and (109) we get following dependence tensor-to-scalar ratio from spectral tilt of the scalar perturbations
$r=\frac{8p\sqrt{\frac{1-n}{1+n}}(1-n)(1-n_{S})}{2(1+n)+p(1-n)}.$ (110)
As one can see, for the case $n=0$ this dependence is reduced to (101).
Further, taking into account the observational constraint $r<0.032$ and restrictions (103)–(104) for the same observational values of spectral tilt $\tilde{n}_{S}=n_{S}=0.9663\pm 0.0041$, from (110) we obtain following constraints on the parametrization constant
$0.35<n<1,$ (111)
corresponding to the verifiable inflationary models with potential (105).
Also, we get the following possible values of the velocity of the scalar perturbations $0.70<c_{S}<$ 0.7 from expression (56). Thus, in the sub-luminal regime of the propagation of scalar perturbations, this model corresponds to the observational constraints (43)–(44). Therefore, for the case of scalar-torsion gravity, inflationary models with power-law potential can be verified by observational constraints on the values of the parameters of cosmological perturbations.
VI.3 Inflationary models with exponential potential
Now, we consider the inflationary model with exponential potential [50, 51, 52]
$\tilde{V}(\phi)=\tilde{V}_{0}\exp(-\alpha\phi),$ (112)
where $\tilde{V}_{0}$ and $\alpha$ are a some positive constants.
From (83)–(84) we obtain
$\tilde{\epsilon}=\tilde{\epsilon}_{\ast}=\frac{\alpha^{2}}{2}>0,$ (113)
$\frac{\tilde{\delta}}{\tilde{\epsilon}}=\frac{\tilde{\delta}_{\ast}}{\tilde{\epsilon}_{\ast}}=1.$ (114)
Since relation (114) doesn’t contain additional constant parameters, we can immediately determine the dependence of the tensor-scalar ratio on the spectral index of scalar perturbations. This dependence will be valid for any e-fold number including $7050<\tilde{N}<$ 70.
Thus, from (71)–(72) and (114) we obtain
$\tilde{r}=8(1-\tilde{n}_{S}).$ (115)
After substituting $\tilde{n}_{S}\simeq 0.97$ into (115) one has $\tilde{r}\simeq 0.24$. Thus, the inflationary model with exponential potential (112) doesn’t satisfy the observational constraint (44) for the case of minimal coupling. For the case of non-minimal coupling, from equations (40)–(41) and (112) we get
$V(\phi)\simeq V_{0}\exp\left(-\frac{\alpha(2+n)}{2(1+n)}\phi\right),$ (116)
$F(\phi)\simeq F_{0}\exp\left(-\frac{n\phi}{2(1+n)}\right),$ (117)
where
$V_{0}=\frac{6(1+n)}{(2-n)}\left(\frac{\tilde{V}_{0}}{3}\right)^{\frac{2+n}{2(1+n)}},$ (118)
$F_{0}=\frac{2(1+n)}{(2-n)}\left(\frac{\tilde{V}_{0}}{3}\right)^{\frac{2+n}{2(1+n)}}.$ (119)
After substituting (116) into (88)–(90) we obtain
$\frac{\delta}{\epsilon}=\frac{\delta_{\ast}}{\epsilon_{\ast}}=1+\frac{n}{2}=const,$ (120)
that coincides with relation (75).
Further, from (66)–(67) and (120) we obtain
$r=8\sqrt{\frac{1-n}{1+n}}\,(1-n_{S}).$ (121)
Thus, for $n_{S}\simeq 0.97$ and $r<0.032$, from (121) we get $0.9650<n<$ 0.965. Also, for $0.9650<n<$ 0.965 corresponding velocity of the scalar perturbations (56) takes the values $0.130<c_{S}<$ 0.13.
Therefore, this inflationary model is verified by observational constraints only for the case of a very strong influence of the non-minimal coupling on the velocity of propagation of the scalar perturbations. For this reason, the inflationary models with power-law potential seem more realistic.
VII Conclusion
We testify the influence of the non-minimal coupling of the scalar field and torsion on the cosmological parameters based on the power-law relationship between the coupling function and the Hubble parameter $F=(H/\lambda)^{n}$. This parametrization implies the possibility of a model-independent assessment of the effect of non-minimal coupling on cosmological parameters. Within the framework of this approach, this influence was parameterized by one constant parameter $n$. Also, we found restrictions on the value of parametrization constant $-1<n<1$. The influence of the non-minimal coupling on cosmological dynamics, scalar field potential, and parameters of cosmological perturbations was assessed by this approach.
It was also shown that cosmological inflationary models implying a linear relationship between the tensor-scalar ratio and the spectral index of scalar perturbations can be verified by observational constraints on the values of cosmological perturbation parameters due to the influence of the non-minimal coupling between the scalar field and torsion. It was also noted that such an influence leads to a sub-luminal regime of the propagation of scalar perturbations and a decrease in the value of the tensor-scalar ratio for the models under consideration.
Further development of the proposed approach is associated with its generalization to the case of the violation of relations $H=h^{1/(1+n)}$ and $\delta/\epsilon=const$. This generalization leads to a wider class of cosmological models including models with a super-luminal velocity of propagation of scalar perturbations and increasing tensor-to-scalar ratio.
Acknowledgements
BM acknowledges Bauman Moscow State Technical University(BMSTU), Moscow for providing the Visiting Professor position during which part of this work has been accomplished.